OpenStudy (anonymous):

A man drove from Rizal to Taytay in a total of 150 km in 3 hours. Along the highways, he averaged 90 km per hour but slowed down along the towns and averaged 30 km per hour. How long did he drove along the highways?

OpenStudy (anonymous):

(d/90)+((150-d)/30)=3 d/90+5-d/30=3 -2d/90=-2 d/90=1 d=90 km

OpenStudy (anonymous):

It's asking for the number of hours he drove on the highways. The choices are: 1, 1.5, 2, 2.5

OpenStudy (anonymous):

1.5 hours =90 min

OpenStudy (anonymous):

x=time travelled at speed 90 y=time travelled at speed 30 x+y=3 90x+30y=150 ------------ solve the above 2 to get x=1 and y=2 he drove along the highways for 1 hour

OpenStudy (anonymous):

@ Noureldin-2022 (d/90)+((150-d)/30) =3 d/90+5-d/30=3 -2d/90 =-2 d/90=1 d=90 km at highways, speed=90km/hr and distance travelled d=90 km so time = distance/speed = (90 km) / (90km/hr) = 1 hr